A jet 3582 mil travels against a jetstream in 6 hours and 4602 mi with the jetstream in the same amount of time. What is the rate of the jet in still air and what is the rate of the jetstream?

Answer by mananth(12270)

A jet travels 3582 mil against a jetstream in 6 hours and 4602 mi with the jetstream in the same amount of time. What is the rate of the jet in still air and what is the rate of the jetstream?
let speed of jet in still air be x
speed of jetstream be y
3582 /x-y = 6
3582 = 6x -6y
597=x-y
4602 / x+y= 6
6x+6y = 4602
x+y= 767

x-y+x+y= 597 + 767
2x=1364
x= 682 mph speed in still air
plug the value of x
682+y=767
y= 767-683
y= 83 mph whch is the speed of the stream